### v2

# All-optical compact setup for generation of collimated multi-MeV proton beams with a ”snail” target

Abstract
The work considers an optical scheme for collimation of high-energy proton beams using $\sim 10^{5}$ T scale magnetic fields induced in a miniature "snail" target by petawatt or multi-petawatt laser irradiation in ps or fs regime. Such magnetic fields are known to be frozen into hot plasma and exist on at least a hundred of picoseconds time-scale, allowing their use for control of charged particle beams. The high values of the magnetic field along with the compact size perfectly match conditions for an all-in-one optical setup, where first, the laser beam accelerates protons, by, e.g. Target Normal Sheath Acceleration (TNSA) mechanism, and second, the closely positioned snail target is driven to guide the proton beam. An important issue is that the laser drivers for both proton acceleration schemes and the magnetic field generation in the considered targets may have the same properties, and even be parts of one splitted beam. Numerical simulations show that the considered setup can be used for efficient collimation of $\simeq 100$ MeV protons. The collimation effect weakly depends on the fine magnetic field structure and can be observed both for a simple magneto-dipole field profile and for a more complex coaxial-like profiles accounting for the intricate structure of electric currents in the interaction region. The obtained results are interesting for the development of intense laser-driven sources of charged particle beams with low divergence and high energy of accelerated particles.
laser plasma, laser-driven magnetic fields, laser-plasma acceleration, magnetic collimation
preprint: AIP/123-QED
I Introduction
With the development of technologies for generation and amplification of ultra-short laser pulses, their intensities far exceeding $10^{18}$ W/cm2 became attainable Yoon et al. (2021). When such intense laser fields interact with matter, they can effectively accelerate electrons Mourou, Tajima, and Bulanov (2006); Pukhov, Sheng, and Meyer-ter Vehn (1999); Pukhov (2002); Rosmej et al. (2020); Esarey, Schroeder, and Leemans (2009) and ions Maksimchuk et al. (2000); Snavely et al. (2000); Roth and Schollmeier (2016). In the latter case, the acceleration nature usually is the charge separation fields formed in the process of laser irradiation of a solid target. Laser-driven sources of high-energy charged particles can offer several advantages over other schemes based on conventional accelerators. They present a very compact setup where strong accelerating fields allow reaching very high acceleration gradients Katsouleas et al. (1997). At the same time, the number of laser-accelerated particles can also be quite high both for electron Rosmej et al. (2020) and ion Roth and Schollmeier (2016) acceleration. Laser-accelerated electron beams carrying high charges and high electric currents can be used for intense gamma-ray and neutron production Günther et al. (2022); Tavana et al. (2023), research in the field of nuclear photonics Ma et al. (2019); Nedorezov, Rykovanov, and Savel’ev (2021). In addition, directed beams of accelerated electrons can be used as a driver for ion acceleration Roth and Schollmeier (2016). Laser-accelerated ions and, in particular, protons, can be used e.g. as a diagnostic tool for probing ultra-fast transient phenomena, like pulsed electromagnetic fields generated with powerful lasers Willingale et al. (2010); Law et al. (2016a); Gao et al. (2016); Palmer et al. (2019); Bradford et al. (2020), and to produce warm dense matter for equation-of-state studies Patel et al. (2003); Dyer et al. (2008). Fast protons are also of a great interest in the context of fast ignition approach to inertial confinement fusion Temporal, Honrubia, and Atzeni (2002); Bychenkov et al. (2001). Also, laser-accelerated protons offer a promising potential for medical applications, i.e. cancer treatment with the hadron therapy Bulanov et al. (2002).
Despite their advantages, laser-driven sources of charged particle beams typically suffer from a poor beam quality and a large divergence in comparison to conventional accelerators, limiting their use for potential applications such as the radiation therapy Fourkal et al. (2002, 2003). It is, however, possible to reduce the beam angular spread by exploiting strong electromagnetic fields for beam collimation. For example, it has been shown Bailly-Grandvaux et al. (2018); Santos et al. (2018) that laser-induced magnetic fields of about $\left(500-600\right)$ T can be used to magnetize a solid target and achieve a guiding of relativistic electron beams generated by irradiation of a target surface with a high-energy picosecond laser pulse. In this case, interaction with the magnetic field in the process of propagation through the magnetized target decreased the electron beam divergence and caused a 5-fold increase of the energy-density flux at the rear edge of a $60$-$\upmu$m-thick target Bailly-Grandvaux et al. (2018); Santos et al. (2018). Strong laser-driven electromagnetic fields have also been applied for ion collimation, see Refs. Kar et al. (2016); Ahmed et al. (2017); Liu et al. (2024). There, travelling electromagnetic pulses were excited in helical coils and the fields associated with those pulses were shown to be capable of a guided post-acceleration of $\sim\left(1-10\right)$ MeV TNSA-produced protons.
A practical point for applications is the simplicity and robustness of a possible setup which produces a collimated proton beam. So called capacitor-coil targets Santos et al. (2015, 2018); Law et al. (2016b); Daido et al. (1986); Courtois et al. (2005) are well studied in ns - kJ regime, where they are able to create kilotesla-scale magnetic fields for several ns. However, in a ps regime, their efficiency seems to be much lower, see, e.g. Ehret et al. (2023). This means that in order to build an optically controlled compact source of collimated protons with use of capacitor-coil targets, probably two kinds of laser channels are required: the one for a high-power short (fs or ps) beam, and the other for a high-energy long (ns or hundreds of ps) beam. There are possible solutions aiming to simplify a possible setup, e.g. the travelling discharge, discussed in e.g. Ahmed et al. (2017), may use a short powerful beam, even the same which creates the TNSA protons.
As an alternative to the setups with electromagnetic pulses in a helical coil, the beam collimation may be achieved with the use of strong quasi-stationary magnetic fields excited in miniature "snail" targets in the picosecond Ehret et al. (2022) or even in a femtosecond Bukharskii and Korneev (2023) regime. This approach requires two laser beams but allows for a possibility of an additional adjustment freedom. With a relatively modest $50$ J of invested laser energy in the form of sub-petawatt $\simeq 0.5$ ps laser pulse, the magnetic field strength in such a target can reach kilotesla level Ehret et al. (2022), while its life-time exceeds $\simeq 100$ ps. With more powerful laser drivers, even higher $\sim 10^{5}$ T magnetic fields are predicted Bukharskii and Korneev (2023).
Although the snail targets are experimentally realized and showed a very good robustness, experimental studies related to application of these targets to proton collimation are yet at the planning stage. Previously, it was found, that depending on the interaction parameters, the structure of the magnetic field in the snail target cavity may be different. However, as the currents are bound to the target surface, at least in the central part the direction of the fields is close to that of the snail axis. Nevertheless, it is not yet clear how much is the internal structure of the field important for the collimation properties of the target and which magnetic field scales are required for collimation in certain proton energetic ranges. The main goal of this work is to consider different possible situations and to demonstrate that an efficient collimation of fast protons in magnetic fields, created by short powerful laser beams in a snail target is achievable with modern and perspective laser facilities. The considered setup for producing high-energy collimated proton beams is illustrated in Fig. 1, (a). The beam is assumed to be produced via TNSA-mechanism Roth and Schollmeier (2016) and propagates through the "snail" microcoil, where it interacts with the collimating magnetic field. The resulting beam profile is analyzed at the distances of several tens of mm behind the snail target.
II Magnetic field structure and parameters
In the simulation presented below, the magnetic field profile was defined for a given electric current geometry using the Biot–Savart law. The current geometry was chosen to reproduce either "uniform" or "coaxial" B-field profiles. By the "uniform" profile we mean hereafter the field profile which is topologically similar to that of the magnetic dipole. The "coaxial" profile was obtained in earlier studies in 2D PIC simulations, when the snail targets are irradiated with very intense but very short laser pulses, see, e.g. in Bukharskii and Korneev (2023). This specific magnetic field structure appears when the current of laser-accelerated electrons is strong enough to magnetize an internal part of the coil. Fig. 1, (b,c) demonstrates the two magnetic field profiles used in this work. However, in many cases, at least for a ps-duration laser driver, the internal volume with an opposite polarity of the magnetic field was not seen in simulations Korneev, D’Humières, and Tikhonchuk (2015); Korneev (2017); Ehret et al. (2022).
The current geometry providing the magnetic field structure is shown in Fig. 1, panels (b,c), with arrows. The discharge current is marked with the light green arrow, the current of deflected laser-accelerated electrons is shown with the light-blue arrow, and the red arrow indicates the laser beam propagation. The target profile is shown with the dark gray mask. Both the uniform and the coaxial magnetic field profiles are shown in the figure. The electric field is of course also excited in the system as the target is being charged positively under the action of the intense driver Ehret et al. (2022), though its effect in the considered situation is weak, see the discussion below.
Figure 1: Sketch of the proposed setup for producing high-energy collimated proton beams (a) and spatial distribution of the transverse magnetic field component $B_{z}$ inside the snail target: (b) a simple uniform B-field profile produced by the discharge current along the inner "snail" surface (light green arrow), (c) coaxial B-field profile reproducing the one obtained in 2D PIC simulations in Bukharskii and Korneev (2023), formed by the discharge current of the bulk electrons (light green arrow) and the current of deflected laser-accelerated electrons (light blue arrow). Direction of the incident laser pulse used to excite the B-field in "snail" microcoil is schematically shown with the red arrow.
The propagation of high-energy proton beams through the considered electromagnetic structures was modeled using a test-particle approach. The simulations were performed using a self-developed code written in the high-level programming language Python Van Rossum and Drake (2009), complemented by the NumPy library Harris et al. (2020) and an open-source JIT-compiler Numba Lam, Pitrou, and Seibert (2015) transforming a regular Python code into a fast machine code and allowing parallelization for multiple CPUs. A beam of protons was sent through the cavity of the "snail" target along the axis ($z$-axis in Fig. 1, (b,c)). For the propagating particles in the given fields, relativistic equations of motion with the Lorentz force were solved numerically for each particle in the beam using the Boris integration scheme Boris (1970). The fields were assumed to be stationary during the time required for a multi-MeV proton to cross the target region ($100$ MeV protons cross $\simeq 100$ $\upmu$m-long high field region in less than $1$ ps, while the magnetic field changes much slower, on the temporal scale of a few hundreds of ps Ehret et al. (2022); Bukharskii and Korneev (2023)). The interaction between different particles in the proton beam was neglected as space-charge effects are typically low for TNSA-produced proton beams Santos et al. (2018). Proton scattering and energy losses due to collisions with plasma formed in the target cavity were also assumed to be negligible for the considered projectile energies exceeding $10$ MeV scales, taking into account small plasma cloud spatial scales of $\sim 100$ $\upmu$m and characteristic density values of $10^{19}$..$10^{21}$ cm-3. The magnetic fields in the target region were calculated on a 3D dimensional grid with 2 $\upmu$m resolution under the assumption that they are produced by the fixed electric currents defined by the contours in Fig. 1, (b,c). For this purpose, each of the assumed current paths, i.e. one in the case of the uniform B-field profile (light green in Fig. 1(b)) or two in the case of the coaxial B-field profile (light green and light blue in Fig. 1(c)), were divided into $N_{I}=100$ infinitely thin linear segments of approximately equal lengths. The resulting magnetic field at position $\mathbf{r}$ was found as a sum of contributions from all the $N_{I}$ segments according to the Biot-Savart law:
$\mathbf{B}\left(\mathbf{r}\right)=\frac{\mu_{0}}{4\pi}\sum_{i=1}^{N_{I}}\frac{I\,d\mathbf{l}_{i}\times\mathbf{r}_{i}^{\prime}}{\lvert\mathbf{r}_{i}^{\prime}\rvert^{3}},$
where $I$ is the current value, $d\mathbf{l}_{i}$ is the vector drawn from the start to the end of the $i$-th linear segment, $\mathbf{r}_{i}^{\prime}=\mathbf{r}-\mathbf{r}_{i}$ is the vector from the position $\mathbf{r}_{i}$ of the $i$-th linear segment to the point $\mathbf{r}$ where the field is calculated, and $\mu_{0}$ is the magnetic constant. The current value was $\approx 7\cdot 10^{5}$ A, which corresponds to the $10^{5}$ T scale magnetic fields obtained in 2D PIC simulations in Bukharskii and Korneev (2023). The electric fields were determined under the assumption that the target is a perfect conductor charged to a certain potential. The latter was set to $200$ kV, similar in scale to the value retrieved from the experimental data in Ehret et al. (2022). Although, it should be noted that for the considered $\simeq 100$ MeV protons electric fields corresponding to potentials of a few hundred kilovolts have a little effect on trajectories of fast protons – further modeling of proton beam propagation showed the resulting beam profile remained almost unchanged if the electric potential was varied in $\left(0-1000\right)$ kV range. Particles were originated from a point source placed $0.5$ mm away from the "snail" center, see the related discussion in Section IV. Their velocities were defined according to their kinetic energy, and their initial angles were set so that the angular distribution of the beam presented a Gaussian profile with the full width at half maximum (FWHM) of $20^{\circ}$, which corresponds to the half opening angle $\theta/2=10^{\circ}$. The beam was directed along the "snail" axis and registered at the "detector" plane placed $50$ mm away from the the target plane. A total of $10^{6}$ particles were used in the simulation to accumulate sufficient statistics in the detector plane.
Figure 2: Results of test-particle simulations of proton beam transport through the "snail" cavity for the uniform $\sim 10^{5}$ T magnetic field and the electric field corresponding to $200$ kV target surface potential: (a) initial beam profile in the detector plane if no fields are imposed on the beam; (b-f) beam profiles in the detector plane modified due to the interaction with the B-field around the target for different projectile energies – $50$ MeV, $80$ MeV, $120$ MeV, $150$ MeV and $200$ MeV, respectively; (g) angular distribution for the divergent proton beam (black curve) and collimated beam of $120$ MeV protons (blue curve).
III Proton beam collimation in the electromagnetic fields of the "snail" target
In the first set of simulations, the proton beam propagation was studied for the uniform B-field profile, see panel (b) in Fig. 1. Projectile energies in range $\left(50-200\right)$ MeV were considered. The obtained results are presented in Fig. 2. The presented profiles show that the imposed magnetic field causes a significant redistribution of the particles in the beam. Initially, the beam has a wide profile that occupies most of the detector area, see Fig. 2, (a). However, when the beam passes through the cavity with the quasi-stationary $10^{5}$ T scale B-field, protons deflect and form three different regions which can be observed in the detected profiles, panels (b-f). They consist of the "external" region, the "void" region with a bright boundary and the "inner" collimated beam region. The first one, i.e. the "external" region, encompasses all of the area far from the center of the detector plane and it stays almost unmodified relative to the original beam profile, as no significant changes of proton flux are observed in it. This is not surprising as this region corresponds to protons that do not pass through the "snail" cavity, but instead go around it. As the magnetic field strength is considerably greater inside the cavity, it has a little effect on these protons and just slightly deflect them. The second region represents a ’void’ with a very low signal. The outer boundary of this region, however, is very bright and has a pronounced caustics. The third and final region is the one inside the ’void’ close to its center. This area corresponds to protons that pass through the center of the target close to its axis. These protons are effectively collimated for almost all of the presented cases, see Fig. 2, (b-f).
The achieved degree of collimation depends on the projectile energy. For the $\sim 10^{5}$ T magnetic field produced by the electric current of $7\cdot 10^{5}$ A shown in Fig. 1, (b), the optimal range of energies for which a high degree of collimation is achieved is around $\left(80-120\right)$ MeV. If the energy is lower, protons undergo very strong deflection and focus at some point before the detector plane, instead of collimating into a quasi-parallel beam, thus making the resultant pattern at the center of the detector plane irregular and prolonged in certain directions. Protons with higher energies, on the other hand, are under-collimated, and the spot they produce is also irregular. Its shape resembles the shape of the "snail" cavity, while its size is significantly higher than that for protons within the optimal energy range. At the same time, the maximum proton flux for protons outside of the optimal energy range is several times lower.
In Fig. 2, (g), angular distributions are presented for the initial divergent beam and the collimated beam of $120$ MeV protons. The presented plot shows the three aforementioned regions the proton beam is divided into: (I) the "external" region extending from about $6^{\circ}$ upwards; (II) the ’void’ region from $1^{\circ}$ to $4^{\circ}$, enclosed by a bright boundary at $\left(4.0^{\circ}\pm 0.5^{\circ}\right)$; and (III) the "inner" collimated spot region at $\left(0^{\circ}-1^{\circ}\right)$. Due to the redistribution of protons under the action of the magnetic field their flux per unit solid angle along the axis of the beam increases by a factor of $\simeq 8$, while the characteristic divergence angle of the beam decreases from the initial $10^{\circ}$ to the value below $0.5^{\circ}$. This corresponds to more than $100$ times decrease of the solid angle covered by the beam.
Figure 3: Results of test-particle simulations of proton beam transport through the "snail" cavity for the coaxial $\sim 10^{5}$ T magnetic field and the electric field corresponding to $200$ kV target surface potential: (a) initial beam profile in the detector plane if no fields are imposed on the beam; (b-f) beam profiles in the detector plane modified due to the interaction with the B-field around the target for different projectile energies – $50$ MeV, $80$ MeV, $110$ MeV, $150$ MeV and $200$ MeV, respectively; (g) angular distribution for the divergent proton beam (black curve) and collimated beam of $110$ MeV protons (red curve).
The obtained collimation properties of the target may be rather qualitatively described in terms of a magnetic lens approach. Following Reiser (2008), in the paraxial approximation, the focal length $f$ for a solenoidal magnetic field lens with the peak magnetic field $B_{0}$ and the effective field length
[…]
is
$f=\frac{l}{\Phi\sin\Phi},$ (1)
[…]
In the second set of simulations, the proton beam propagates through the target with the coaxial-shape $10^{5}$ T scale magnetic field, see Fig. 1(c). The obtained proton beam profiles in the "detector" plane are shown in Fig. 3, where panels (b-f) correspond to different proton energies in the range $\left(50-200\right)$ MeV. As can be seen, at the optimal and sub-optimal energies, i.e. $\left(50-120\right)$ MeV, the profiles do not differ qualitatively from the case of the uniform B-field, see Fig. 3(b-d), while for higher proton energies the coaxial B-field profile presents some more intricate structures in the proton patterns, which are shown in Fig. 3(e),(f). The optimal proton energy for which the collimation effect is the highest in the case of the coaxial B-field profile is slightly lower than that for the uniform B-field distribution ($120$ MeV) and equals to $110$ MeV. The shape of the central spot, corresponding to the collimated part of the beam, is very similar, although the peak increase of the proton flux at the center of this spot is somewhat lower and equals to $\simeq 5.5$. The characteristic divergence half-angle of the collimated beam is $\approx 0.5^{\circ}$, similar to that for the uniform B-field distribution.
Figure 4: Comparison of the results of test-particle simulations of proton beam transport through the "snail" cavity for different B-field profiles: (a,b) beam profiles in the detector plane for the uniform and coaxial B-field profiles, with projectile energies of $120$ MeV and $110$ MeV, respectively; (c) angular distribution for the beam of $120$ MeV protons collimated by the uniform B-field (blue curve) and $110$ MeV protons collimated by the coaxial B-field (red curve).
Fig. 4 shows an explicit comparison of the resulting proton beam profiles for the uniform and coaxial collimating field structures. Panels (a,b), where the two-dimensional distributions are presented, illustrate that the coaxial field structure is slightly inferior to the uniform B-field distribution of the same strength – with the same color scale, the proton beam intensity in the focal spot appears to be less bright. The spot itself is also somewhat larger in panel (b), and proton density does not drop as steeply at its edges as for the uniform B-field case shown in panel (a). The analysis of the angular distributions also show that less particles propagate along the target axis during interaction with the coaxial $10^{5}$ T B-field in comparison to the uniform $10^{5}$ T B-field profile. However, the difference amounts to only a few tens of percent, and thus, both the coaxial and the uniform B-field profiles in the considered range of parameters can be used effectively for the proton beam collimation.
IV Discussion
The calculations described in Sec. III were performed under the assumption that the magnetic field induced around the target remains static and thus does not excite vortex electric field. To justify the validity of such assumption for the considered problem, an additional set of simulations was performed to study the effect of the inductive electric field on particle collimation. This field was calculated as
$\mathbf{E_{i}}=-\frac{1}{c}\frac{\partial\mathbf{A}}{\partial t},$ (2)
where $\mathbf{A}$ denotes the retarded vector potential:
$\mathbf{A}\left(\mathbf{r},t\right)=\frac{1}{c}\int_{\Omega}\frac{\mathbf{J}\left(\mathbf{r^{\prime}},t^{\prime}\right)}{\left|\mathbf{r}-\mathbf{r^{\prime}}\right|}d^{3}\mathbf{r^{\prime}}.$ (3)
The current in Eq. (3) was constant along each of the selected paths, see Fig. 1, panels (b,c), and decayed linearly in time from $\sim 7\cdot 10^{5}$ A (corresponding magnetic field is $B\sim 10^{5}$ T), down to zero. To demonstrate the effect, different decay times were considered, $\tau_{d}=100$ ps, $\tau_{d}=10$ ps and $\tau_{d}=1$ ps, where the value $\tau_{d}=100$ ps corresponds to the one retrieved from experimental data for "snail" targets in Ehret et al. (2022). Resulting beam profiles are shown in Fig. 5.
Figure 5: Results of test-particle simulations of proton beam transport through the "snail" cavity for different B-field profiles, uniform (top row) and coaxial (bottom row), with the inductive electric field caused by the time-varying magnetic field taken into account. Panels (a,d) correspond to the magnetic field decay time $\tau_{d}=100$ ps, panels (b,e) – to $\tau_{d}=10$ ps and panels (c,f) – to $\tau_{d}=1$ ps.
The inductive electric field induced by the decaying annular current is also annular in the target cavity region. Its main component is $E_{\phi}$ where $\phi$ denotes the azimuthal angle measured in the plane perpendicular to the beam propagation axis. Such electric field changes projectile’s azimuthal velocity component, which in turn causes a change of the radial component velocity component, as $\mathbf{v}_{\phi}\times\mathbf{B}$ force is directed radially inside the "snail" cavity. From the obtained distributions for the uniform B-field profile it follows that if the B-field changes on the time-scale $\sim\left(10-100\right)$ ps, the described effect is insignificant, see panels (a) and (b) in Fig. 5, where the beam profiles change only very slightly in comparison to the those presented in Sec. III, with a decrease of peak proton density at the center of the collimated beam not exceeding a few percent. For $\tau_{d}=1$ ps, however, the inductive electric field strength appears to be strong enough to cause a significant redistribution of protons and reduce the collimation efficiency. This case is shown in panel (c) of Fig. 5. The same applies to the beam profiles obtained for the coaxial B-field structure, shown in panels (d-f). As for the uniform B-field profile, the inductive electric field impairs collimation only for $\tau_{d}=1$ ps, see panel (f) in Fig. 5. Therefore, for the significant inductive electric field contribution in the considered scheme, the magnetic field has to change on $\sim 1$ ps time-scale, which is much faster than the B-field decay rate measured experimentally Ehret et al. (2022) and observed in simulations Bukharskii and Korneev (2023). Thus, in the considered range of parameters, the inductive fields do not play a significant role in the beam collimation. The static electric fields for realistic target potentials up to $1$ MV also appear to be weak to cause a significant deflection of $\simeq 100$ MeV protons, as it was verified by an auxiliary set of simulations. Thus, we can conclude that for the considered high energies of protons and strong values of the magnetic field, the collimation effect is mainly defined by the quasistationary magnetic field induced around the target.
Another important point worth discussing is the used model of the proton beam – in Sec. III a simple ideal case was considered, where a monoenergetic proton beam is emitted from a point-like source, while in reality the source has finite dimensions and some energy spread. Assuming that the proton beam is produced via the TNSA mechanism during an ultrashort laser pulse irradiation of a thin metallic foil. The imaging properties of such a source, despite relatively large transverse size of the emitting region of $\sim 100-200$ $\upmu$m, can be reproduced by assuming a much smaller source shifted by $\sim 100$ $\upmu$m from the foil surface Borghesi et al. (2004). The effective size of such a source can be as small as $\simeq 5-10$ $\upmu$Borghesi et al. (2004); Kugland et al. (2012). As discussed in Kugland et al. (2012), finite source size may lead to blurring of proton images, i.e. beam profiles, and can be neglected if the source size $d_{s}$ is much lower than the probed field size $d$: $d_{s}\ll d$. Here, assuming $d_{s}\simeq 5-10$ $\upmu$m the condition $d_{s}\ll d$ is satisfied as $d\simeq 50$ $\upmu$m, and blurring related to the finite source size should be insignificant, which was verified by an auxiliary set of simulations. There, the initial transverse proton positions were chosen randomly inside circles with different diameters, $d_{s}=0.2$ $\upmu$m, which corresponds to an almost point-like source, as well as $d_{s}=5$ $\upmu$m and $d_{s}=10$ $\upmu$m which describe finite-size sources. Results of the performed simulations for the uniform B-field profile are shown in Fig. 6. As can be seen, blurring becomes significant only for $d_{s}=10$ $\upmu$m, see Fig. 6(c). In this case, the peak proton density drops by $\approx 35$ in comparison to the point-like source case, shown in panel (a) of the same figure. For $d_{s}=5$ $\upmu$m, the peak proton density decrease due to the blurring amounts to a modest $10$. Similar results were obtained assuming the coaxial B-field profile. Thus, for the considered scheme the point source assumption remains viable for $d_{s}\lesssim 5$ $\upmu$m as the transverse source size in this case is much smaller than the collimating field transverse size, and the collimation efficiency remains almost unaffected.
Figure 6: Results of test-particle simulations of proton beam transport through the "snail" cavity for the uniform B-field, $120$ MeV proton energy and different source sizes $d_{s}$: (a) point-like source with $d_{s}=0.2$ $\upmu$m, (b,c) finite-size sources with $d_{s}=5$ $\upmu$m and $d_{s}=10$ $\upmu$m, respectively.
In terms of the energy spectrum, TNSA-produced protons are usually characterized by a wide energy spread Roth and Schollmeier (2016). Although, there have been a number of reports of quasi-monoenergetic ion beams produced via TNSA-mechanism using specifically-designed and prepared targets Hegelich et al. (2006); Schwoerer et al. (2006); Ter-Avetisyan et al. (2006) – there, the energy spread of accelerated ions did not exceed $20$. Despite relative difficulties in obtaining monoenergetic ion beams with TNSA, it should be mentioned that for practical applications involving interaction of the TNSA-produced proton beams with thick targets the target itself plays a role on an energy filter due to the presence of well-localized region of ion losses in matter, the so-called Bragg peak. In this case, only a small fraction of the whole beam energy from a certain part of the spectrum, is deposited in the target at a certain depth. This allows for the used description of the monoenergetic proton beams in Sec. III. Nevertheless, it is worth studying how the energy spread effects the collimation efficiency. The study was performed both for the uniform and coaxial B-field profiles, assuming three different values of the energy spread – $\delta E_{0}=\pm 1$, i.e. almost a monoenergetic beam, as well as $\delta E_{0}=\pm 20$ and $\delta E_{0}=\pm 50$. Results are summarized in Fig. 7, panels (a-f).
Figure 7: Results of test-particle simulations of a non-monoenergetic proton beam transport through the "snail" cavity: (a-c) $120$ MeV protons in the uniform B-field with different values of energy spread, $\delta E_{0}=1$, $\delta E_{0}=20$ and $\delta E_{0}=50$, respectively; (d-f) $110$ MeV protons in the coaxial B-field with the same values of the energy spread, $\delta E_{0}=1$, $\delta E_{0}=20$ and $\delta E_{0}=50$, respectively; (g-h) beam profiles for the uniform and coaxial B-field profiles assuming an exemplary angular-energy distribution for the TNSA-mechanism.
From the obtained distributions it follows that the decrease of the collimation efficiency is relatively modest, even for $\delta E_{0}=50$ – the collimated spot retains almost the same shape as in the case of a quasi-monoenergetic beam shown in Fig. 7(a,d), but the peak proton density at the center of this spot slightly drops. For $\delta E_{0}=20$ which has already been successfully achieved in several experiments on laser-driven ion acceleration Hegelich et al. (2006); Schwoerer et al. (2006); Ter-Avetisyan et al. (2006), this drop amounts to only $\simeq\left(6-12\right)$, see Fig. 7(b,e). For higher $\delta E_{0}$ values, the decrease of the peak proton density depends on the B-field structure – for the uniform profile, it still remains relatively small, about $12$, see Fig. 7(c), while for the coaxial profile it amounts to $\approx 30$, see Fig. 7(f).
[…]
Finally, it is also worth discussing how the beam profile changes as it propagates along its axis. Above, distributions in a single plane at a fixed distance of $50$ mm from the target plane were provided. To illustrate the beam evolution as is propagates in the longitudinal direction, transverse beam profiles were compared at different distances between the detector and the target plane – $40$ mm, $50$ mm and $60$ mm (i.e. in ±$\pm 10$ mm distance range from the plane where the profiles had been calculated previously) and additionally at $250$ mm (i.e. $5$ times farther from the target than before). The results are shown in Fig. 8.
Figure 8: Results of test-particle simulations of proton beam transport through the magnetized "snail" target: (a-c) beam profiles for $120$ MeV protons and for the uniform B-field at different distances from the target plane – $40$ mm, $50$ mm and $60$ mm, respectively; (d-f) beam profiles for $110$ MeV protons and the coaxial B-field at the same three distances from the target plane – $40$ mm, $50$ mm and $60$ mm, respectively; (g) the beam profile for $120$ MeV protons and the uniform B-field at $250$ mm distance from the target plane; (h) the beam profile for $110$ MeV protons and the coaxial B-field at $250$ mm distance from the target plane.
Examination of the subsequent transverse cuts at $\left(50\pm 10\right)$ mm both for the uniform (Fig. 8(a-c)) and the coaxial B-field profiles (Fig. 8(d-f)) indicates that the collimated beam in not completely parallel and retains some residual divergence, as was already illustrated, e.g. in panels (c) of Figs. 2 and 3. This results in a linear decrease of the central spot size with distance, so that at the $250$ mm distance it appears to be about $5$ times larger than at the $50$ mm distance, while the peak proton concentration drops as the distance squared, so that $250$ mm away from the target it is about twenty times lower than at the $50$ mm distance, see Fig. 8(a-f) and Fig. 8(g,h). This can be in part a consequence of an imperfectly symmetric field structure, i.e. asymmetry of the field distribution relative to the beam propagation axis. Yet, the imposed magnetic field has a strong positive effect on the peak proton density – in the central spot corresponding to protons that pass through the target cavity for all distances it increases by $\simeq\left(7-14\right)$ times, depending on the field structure. At $250$ mm distance from the target, the inner structure of the central spot visible in panels (a-f) becomes resolved, see panels (g,h) of Fig. 8. Interestingly, it may also have additional, "inner" caustics, which are better visible in the case of the uniform field, see panel (g). A small spot formed by collimated protons is visible close to the center of the profile in both panels (g) and (h). At a large distance it can be seen that this most collimated proton fraction slightly changes its propagation direction – the spot is shifted from $\left(0,0\right)$ position at the detector plane, by the shift is small, with the deflection angle relative to the propagation axis of $\simeq\left(0.3^{\circ}-0.4^{\circ}\right)$.
V Conclusion
In this work, collimation of $\simeq 100$ MeV protons by electromagnetic structures with $10^{5}$ T scale magnetic fields induced inside a "snail" target was studied using test-particle simulations. Two qualitatively different magnetic field profiles were considered – a simple uniform distribution formed solely by the discharge current flowing along the inner "snail" surface, and a more complex coaxial structure, where in addition to the discharge current, currents of laser-accelerated electrons deflected in the forming magnetic field also provide a significant contribution to the resulting B-field profile. The both considered B-field profiles were shown to be suitable for collimation of high-energy protons. The uniform structure, according to the results of the presented numerical modeling, can be used to collimate $120$ MeV protons providing $8$ times increase of the proton flux at the center of the collimated beam. The coaxial B-field profile was shown to be slightly inferior to the uniform structure – according to the performed simulations, it can be used for collimation of $110$ MeV protons providing $5.5$ times increase of their concentration. Despite about a $30$ difference in the maximum achieved proton flux, both the uniform and the coaxial B-field profiles were shown to decrease the beam divergence by more than a factor of $10$, from $10^{\circ}$ to $\approx 0.5^{\circ}$ (FWHM). This corresponds to $\geq 100$-fold decrease of the solid angle covered by the proton beam.
The collimation effect for the scheme considered here is attributed mainly to the action of the $\sim 10^{5}$ T scale magnetic field, while static and inductive electric fields for realistic parameters do not have a significant impact on the collimation efficiency. The effect of finite source size appears to be negligible, as long as it stays at least an order of magnitude lower than the size of the collimating cavity. Here, in order not to impair the collimation efficiency, the effective source size has to be $\lesssim 5$ $\upmu$m. Although, the considered system can be used to collimate TNSA-produced beams with a wide energy spectrum, it is better suited for the collimation of quasi-monoenergetic beams with energy spread not exceeding $\simeq 20$. However, collimation of beams with wide energy spectrum can be improved by fine-tuning of the B-field decay time – it can be optimized in such a way, that the field’s decay rate matches the time of flight for protons with a particular energy to provide their efficient collimation, turning the considered system into an achromatic magnetic lens. In this case, protons with different energies are assumed to approach the magnetic field region in different times, the lower the energy the greater this time, depending on the initial proton source position. For slower protons the magnetic field needed for their efficient collimation is lesser, and the suitable values may be in principle adjusted using different materials, thicknesses or structures of the targets.
It was found that although, the beam retains some residual divergence $\simeq 1^{\circ}$, probably as a consequence of its non-axisymmetric structure, the magnetic field of the snail target still greatly increases proton concentration at the center of the detector plane – a $\simeq\left(7-14\right)$ increase is observed at all distances from the target plate relative to the initial divergent beam case. The proposed method can be used at modern and perspective petawatt and multi-petawatt laser facilities (e.g. Danson et al. (2019); Kostyukov et al. (2023)) as a part of an integrated all-optical acceleration and guiding setup. In such a scheme, e.g. the TNSA-produced protons created by one laser pulse may pass the optically-driven electromagnetic structure of the considered "snail"-type target. Such a source of fast protons is expected to have an exceptional brightness, low divergence and high energy of accelerated particles, and it would be of interest for fundamental studies and prospective applications.
VI Acknowledgements
The work was funded by the Russian Science Foundation under Grant No. 24-22-00402. The authors acknowledge the NRNU MEPhI High-Performance Computing Center and the Joint Supercomputer Center of RAS.